# On zeros of polynomials associated with Heun class equations

### v2

@@ 4. Perturbative expansions for zeros
The expansion of the zero of $c_{m+1}(B)=0$ labelled by $-D_{k}$ for $k=2,\cdots,m-2$ up to the term $s^{2}$ is written as
(4.38) $B_{k}^{(m+1)}=-D_{k}+\Bigl{[}-E_{k}-\frac{k(k-1+\gamma)}{D_{k}-D_{k-1}}F_{k}-\frac{(k+1)(k+\gamma)}{D_{k}-D_{k+1}}F_{k+1}\Bigr{]}s$
+[$\quad+\Bigl{[}\Bigl{\{}\frac{k(k-1+\gamma)F_{k}}{(D_{k}-D_{k-1})^{2}}+\frac{(k+1)(k+\gamma)F_{k+1}}{(D_{k}-D_{k+1})^{2}}\Bigr{\}}$
@@
$\quad\quad-\frac{k(k-1+\gamma)F_{k}}{(D_{k}-D_{k-1})^{2}}\Bigl{\{}E_{k-1}+\frac{(k-1)(k-2+\gamma)F_{k-1}}{D_{k}-D_{k-2}}\Bigr{\}}$
$\quad\quad-\frac{(k+1)(k+\gamma)F_{k+1}}{(D_{k}-D_{k+1})^{2}}\Bigl{\{}E_{k+1}+\frac{(k+2)(k+1+\gamma)F_{k+2}}{D_{k}-D_{k+2}}\Bigr{\}}\Bigr{]}s^{2},$
and for $k=0,1$ is written as
(4.39) $B_{1}^{(m+1)}=-D_{1}+\Bigl{[}-E_{1}-\frac{\gamma}{D_{1}-D_{0}}F_{1}-\frac{2(1+\gamma)}{D_{1}-D_{2}}F_{2}\Bigr{]}s$
+[$\quad+\Bigl{[}\Bigl{\{}\frac{\gamma F_{1}}{(D_{1}-D_{0})^{2}}+\frac{2(1+\gamma)F_{2}}{(D_{1}-D_{2})^{2}}\Bigr{\}}\Bigl{\{}E_{1}+\frac{\gamma F_{1}}{D_{1}-D_{0}}+\frac{2(1+\gamma)F_{2}}{D_{1}-D_{2}}\Bigr{\}}$
(4.40) $B_{0}^{(m+1)}=-D_{0}+\Bigl{[}-E_{0}-\frac{\gamma}{D_{0}-D_{1}}F_{1}\Bigr{]}s$
@@ 5. Lamé equation and numerical calculation
It follows from Theorem 4.9 that a great part of the zeros of $c_{m+1}(B)=0$ are expanded as
@@
We try to investigate the zeros of $c_{m}(B)$ numerically by setting the parameters to the particular values, and compare with the results by the approximation from the expansion in equation (5.5). We set
@@
A great part of the zeros of $c_{m+1}(B)=0$ are expanded as
@@
for $k=2,\cdots,m-2$. The zeroth order approximation is $-k^{2}$, the first order approximation is $-k^{2}+(k^{2}/2-3/4)s$, and the second order approximation is $-k^{2}+(k^{2}/2-3/4)s+\{3k^{2}/32-3/32-9/(32(4k^{2}-1))\}s^{2}$. The zeros corresponding to $k=0,1$ are written as
@@
On the other hand, we calculate the zeros of $c_{m}(B)$ numerically by fixing the value $s$ to the particular value. We discuss the case $s=1/100$, which is equivalent to $t=100$. The polynomial $c_{4}(B)$ is written as
@@
The zeros of $c_{40}(B)$ are calculated similarly, and more than 20 zeros of $c_{40}(B)$ coincide with those of the zeros of $c_{30}(B)$ numerically up to 10 digits. Therefore, we can expect that each zero of $d_{2}(B)$ in equation (3.8) is a limit of a zero of $c_{m}(B)$ as $m\to\infty$. We summarize the approximations of the zeros with respect to the expansions on the parameter $s(=1/100)$ and the numerical calculations of the zeros as the following table.
00th approx. $1$st approx. $2$nd approx. zero of $c_{30}(B)$
$k=0$ 00 .$-.0075000000$$-.0074812500$$-.007481156136$
$k=1$ $-1$ $-1.002500000$ $-1.002518750$ $-1.002518844$
$k=2$ $-4$ $-3.987500000$ $-3.987473750$ $-3.987473618$
$k=3$ $-9$ $-8.962500000$ $-8.962425804$ $-8.962425430$
@@
We may expect that each zero of $c_{m}(B)$ converges as $m\to\infty$, although the convergence speed is worse than the case $s=1/100$. We summarize the approximations with respect to the expansions on the parameter $s(=1/2)$ and the numerical calculation as the following table.
00th approx. $1$st approx. $2$nd approx. zero of $c_{40}(B)$
$k=0$ 00 .$-.3750000000$$-.3281250000$$-.3169872981$
$k=1$ $-1$ $-1.125000000$ $-1.171875000$ $-1.183012702$
$k=2$ $-4$ $-3.375000000$ $-3.309375000$ $-3.284829947$
$k=3$ $-9$ $-7.125000000$ $-6.939508929$ $-6.869999689$
@@ 6.1. Mathieu equation
We discuss the reduced singly confluent Heun equation. The recursive relation for the polynomial $c_{m}(B)$ is written in equation (3.21). It follows from Theorem 4.9 that a great part of the zeros of $c_{m+1}(B)=0$ are expanded as
@@
+$\quad+\Bigl{\{}-\frac{1}{8}+\frac{3}{4}\frac{(\gamma-1)^{2}+(\delta-1)^{2}}{(2k-2+\gamma+\delta)(2k+\gamma+\delta)}-\frac{5}{8}\frac{(\gamma-\delta)^{2}(\gamma+\delta-2)^{2}}{(2k-2+\gamma+\delta)^{2}(2k+\gamma+\delta)^{2}}$
@@
We try to investigate the zeros of $c_{m}(B)$ numerically by setting the parameters to the particular values, and compare with the results by the approximation from the expansion in equation (6.5). We set
@@
A great part of the zeros of $c_{m+1}(B)=0$ are written as
@@
We calculate the zeros of $c_{m}(B)$ numerically by fixing the value $s$ to the particular value. We discuss the case $s=2$. The zeros of $c_{8}(B)$ are calculated numerically as
@@
The zeros of $c_{40}(B)$ are calculated similarly, and more than 20 zeros of $c_{40}(B)$ coincide with those of the zeros of $c_{30}(B)$ numerically up to 10 digits. We can expect that each zero of $c_{m}(B)$ converges as $m\to\infty$ for the case $s=2$. We summarize the approximations of the zeros with respect to the expansions on the parameter $s(=2)$ and the numerical calculations of the zeros as the following table.
00th approx. $1$st approx. $2$nd approx. zero of $c_{30}(B)$
$k=0$ 00 $1$ $1.500000000$ $1.378489221$
$k=1$ $-1$ 00 .$-.4166666667$$-.2931662833$
$k=2$ $-4$ $-3$ $-3.033333333$ $-3.035300946$
$k=3$ $-9$ $-8$ $-8.014285714$ $-8.014303906$
$k=4$ $-16$ $-15$ $-15.00793651$ $-15.00793924$
$k=5$ $-25$ $-24$ $-24.00505051$ $-24.00505119$
The $2$nd order approximation is not good for small $k$, but it seems that the approximation would be better if $k$ is getting large.
@@
We can also expect that each zero of $c_{m}(B)$ converges as $m\to\infty$ for the case $s=2i$. We summarize the approximations of the zeros with respect to the expansions on the parameter $s(=2i)$ and the numerical calculations of the zeros as the following table.
00th approx. $1$st approx. $2$nd approx. zero of $c_{30}(B)$
[…]
$k=1$ $-1$ $-1+i$$-.5833333333+i- .$$-.5406395812+1.466873304i- .$
$k=2$ $-4$ $-4+i$ $-3.966666667+i$ $-3.968701175+i$
$k=3$ $-9$ $-9+i$ $-8.985714286+i$ $-8.985730155+i$
$k=4$ $-16$ $-16+i$ $-15.99206349+i$ $-15.99206621+i$
$k=5$ $-25$ $-25+i$ $-24.99494949+i$ $-24.99495018+i$
@@
The real parts of the first two numerical zeros for $c_{30}(B)$ coincide up to 10 digits, although the imaginary parts are $i$ for $k=0$ and $k=1$ on the $2$nd order approximation.
@@
More than 20 zeros of $c_{40}(B)$ coincide with those of the zeros of $c_{30}(B)$ numerically up to 10 digits. We can also expect that each zero of $c_{m}(B)$ converges as $m\to\infty$ for the case $s=i$. We summarize the approximations of the zeros with respect to the expansions on the parameter $s(=i)$ and the numerical calculations of the zeros as the following table.
00th approx. $1$st approx. $2$nd approx. zero of $c_{30}(B)$
[…]
@@ 6.2. Whittaker-Hill equation
We discuss the singly confluent Heun equation. The recursive relation for the polynomial $c_{m}(B)$ is written in equations (3.15) with equation (3.16). It follows from Theorem 4.9 that the expansion of the zero of $c_{m+1}(B)=0$ is written as
@@
+$\quad+\Bigl{\{}-\frac{1}{32}+\frac{P_{6}(k)}{(2k-3+\gamma+\delta)(2k-2+\gamma+\delta)^{3}(2k+\gamma+\delta)^{3}(2k+1+\gamma+\delta)}\Bigr{\}}s^{2}$
@@
We try to investigate the zeros of $c_{m+1}(B)$ numerically by setting the parameters to the particular values, and compare with the results by the approximation from the expansion in $s$. We set
@@
The expansion of the zero of $c_{m+1}(B)=0$ is written as
@@
We calculate the zeros of $c_{m}(B)$ numerically by fixing $s$ and $\alpha$ to the particular values.
We discuss the case $\alpha=5$ and $s=-1/100$. The zeros of $c_{8}(B)$ is calculated numerically as
.$-.02475005544,\;-1.025212444,\;-4.025020000,\;-9.025010357,$
@@
More than 20 zeros of $c_{40}(B)$ coincides with those of the zeros of $c_{30}(B)$ numerically up to 10 digits. Therefore, we can expect that each zero of $d_{2}(B)$ in equation (3.8) is a limit of a zero of $c_{m}(B)$ as $m\to\infty$.
We summarize the approximations of the zeros with respect to the expansions on the parameter $s(=2)$ and the numerical calculations of the zeros as the following table.
00th approx. $1$st approx. $2$nd approx. zero of $c_{30}(B)$
$k=0$ 00 .$-.0250000000$$-.0247500000$$-.02475005544$
$k=1$ $-1$ $-1.025000000$ $-1.025212500$ $-1.025212444$
$k=2$ $-4$ $-4.025000000$ $-4.025020000$ $-4.025020000$
$k=3$ $-9$ $-9.025000000$ $-9.025010357$ $-9.025010357$
$k=4$ $-16$ $-16.02500000$ $-16.02500714$ $-16.02500714$
$k=5$ $-25$ $-25.02500000$ $-25.02500568$ $-25.02500568$
@@
We discuss the case $\alpha=5$ and $s=-20$. A part of the zeros of $c_{16}(B)$ is calculated numerically as
2.051282315, 31.28075111, $2.051282315,\;31.28075111,\;54.63259197-14.18872668i,\;\dots,$
@@
2.051136527, 31.66623015, 50.75984059, $2.051136527,\;31.66623015,\;50.75984059,\;59.89047866-28.89267645i,\;\dots,$
and other zeros are not real numbers. We might expect that $c_{m}(B)$ has a zero around $2.051$ as $m\to\infty$, but it would not be true. It is shown by the numerical calculation that all zeros of $c_{50}(B)$ are not real numbers. A part of the zeros of $c_{89}(B)$ is calculated numerically as
@@
26 zeros of $c_{100}(B)$ are real numbers. We expect that the polynomial $c_{m}(B)$ has zeros around $-11.72870190,\;-37.26280325,\;-59.78916204,\dots$, if $m$ is sufficiently large.

### v3

@@ 4. Perturbative expansions for zeros
On the solution to $c_{m+1}(B)=0$ expressed as
(4.39) $B=-D_{k}-D_{k}^{[1],(m+1)}s-D_{k}^{[2],(m+1)}s^{2}+O(s^{2}),$
we have
(4.40) $D_{k}^{[2],(m+1)}=-\Bigl\{\frac{k(k-1+\gamma)F_{k}}{(D_{k}-D_{k-1})^{2}}+\frac{(k+1)(k+\gamma)F_{k+1}}{(D_{k}-D_{k+1})^{2}}\Bigr\}$
@@
$\quad\quad+\frac{k(k-1+\gamma)F_{k}}{(D_{k}-D_{k-1})^{2}}\Bigl\{E_{k-1}+\frac{(k-1)(k-2+\gamma)F_{k-1}}{D_{k}-D_{k-2}}\Bigr\}$
$\quad\quad+\frac{(k+1)(k+\gamma)F_{k+1}}{(D_{k}-D_{k+1})^{2}}\Bigl\{E_{k+1}+\frac{(k+2)(k+1+\gamma)F_{k+2}}{D_{k}-D_{k+2}}\Bigr\}$
for $k=2,\dots,m-2$, and
(4.41) $D_{1}^{[2],(m+1)}=-\Bigl\{\frac{\gamma F_{1}}{(D_{1}-D_{0})^{2}}+\frac{2(1+\gamma)F_{2}}{(D_{1}-D_{2})^{2}}\Bigr\}\Bigl\{E_{1}+\frac{\gamma F_{1}}{D_{1}-D_{0}}+\frac{2(1+\gamma)F_{2}}{D_{1}-D_{2}}\Bigr\}$
$\qquad\qquad+\frac{\gamma E_{0}F_{1}}{(D_{1}-D_{0})^{2}}+\frac{2(1+\gamma)F_{2}}{(D_{1}-D_{2})^{2}}\Bigl\{E_{2}+\frac{3(2+\gamma)F_{3}}{D_{1}-D_{3}}\Bigr\},$
(4.42) $D_{0}^{[2],(m+1)}=\frac{\gamma F_{1}}{(D_{0}-D_{1})^{2}}\Bigl[E_{1}-E_{0}-\frac{\gamma F_{1}}{D_{0}-D_{1}}+\frac{2(1+\gamma)F_{2}}{D_{0}-D_{2}}\Bigr].$
Note that the expression of $D_{k}^{[1],(m+1)}$ was described in Theorem 4.6.
@@ 5. Lamé equation and numerical calculation
It follows from Theorem 4.9 that a large part of the zeros of $c_{m+1}(B)$ are expanded as
@@
We try to investigate the zeros of $c_{m+1}(B)$ numerically by setting the parameters to the particular values, and compare with the results by the approximation from the expansion in equation (5.5). We set
@@
A large part of the zeros of $c_{m+1}(B)$ are expanded as
@@
for $k=2,\cdots,m-2$. Hence, the zeroth order approximation is $k^{2}$, the first order approximation is $k^{2}+(k^{2}/2-3/4)s$, and the second order approximation is $k^{2}+(k^{2}/2-3/4)s+\{3k^{2}/32-3/32-9/(32(4k^{2}-1))\}s^{2}$. The zeros corresponding to $k=0,1$ are written as
@@
On the other hand, we calculate the zeros of $c_{m+1}(B)$ numerically by fixing the value $s$ to the particular value. We discuss the case $s=1/100$, which is equivalent to $t=100$. The polynomial $c_{4}(B)$ is written as
@@
The zeros of $c_{40}(B)$ are calculated similarly, and more than 20 zeros of $c_{40}(B)$ coincide with those of the zeros of $c_{30}(B)$ numerically up to 10 digits. Therefore, we can expect that each zero of $d_{2}(B)$ in equation (3.8) is a limit of a zero of $c_{m+1}(B)$ as $m\to\infty$. We summarize the approximations of the zeros with respect to the expansions on the parameter $s(=1/100)$ and the numerical calculations of the zeros as the following table.
0th approx. 11st approx. 22nd approx. zero of $c_{30}(B)$
$k=0$ 0 .0075000000-.0075000000 .0074812500-.0074812500 .007481156136-.007481156136
$k=1$ 1-1 1.002500000-1.002500000 1.002518750-1.002518750 1.002518844-1.002518844
$k=2$ 4-4 3.987500000-3.987500000 3.987473750-3.987473750 3.987473618-3.987473618
$k=3$ 9-9 8.962500000-8.962500000 8.962425804-8.962425804 8.962425430-8.962425430
@@
We may expect that each zero of $c_{m+1}(B)$ converges as $m\to\infty$, although the convergence speed is worse than the case $s=1/100$. We summarize the approximations with respect to the expansions on the parameter $s(=1/2)$ and the numerical calculation as the following table.
0th approx. 11st approx. 22nd approx. zero of $c_{40}(B)$
$k=0$ 0 .3750000000-.3750000000 .3281250000-.3281250000 .3169872981-.3169872981
$k=1$ 1-1 1.125000000-1.125000000 1.171875000-1.171875000 1.183012702-1.183012702
$k=2$ 4-4 3.375000000-3.375000000 3.309375000-3.309375000 3.284829947-3.284829947
$k=3$ 9-9 7.125000000-7.125000000 6.939508929-6.939508929 6.869999689-6.869999689
@@ 6.1. Mathieu equation
We describe a result in Section 4 on the reduced singly confluent Heun equation. The recursive relation for the polynomial $c_{m+1}(B)$ was written in equation (3.16). It follows from Theorem 4.9 that a large part of the zeros of $c_{m+1}(B)$ are expanded as
@@
$\quad+\Bigl\{-\frac{1}{8}+\frac{3}{4}\frac{(\gamma-1)^{2}+(\delta-1)^{2}}{(2k-2+\gamma+\delta)(2k+\gamma+\delta)}-\frac{5}{8}\frac{(\gamma-\delta)^{2}(\gamma+\delta-2)^{2}}{(2k-2+\gamma+\delta)^{2}(2k+\gamma+\delta)^{2}}$
@@
We try to investigate the zeros of $c_{m+1}(B)$ numerically by setting the parameters to the particular values, and compare with the results by the approximation from the expansion in equation (6.5). We restrict to the case
@@
In this case, a large part of the zeros of $c_{m+1}(B)$ are written as
@@
We calculate the zeros of $c_{m+1}(B)$ numerically by fixing the value $s$ to the particular value. We discuss the case $s=2$. The zeros of $c_{8}(B)$ are calculated numerically as
@@
The zeros of $c_{40}(B)$ are calculated similarly, and more than 20 zeros of $c_{40}(B)$ coincide with those of the zeros of $c_{30}(B)$ numerically up to 10 digits. Therefore, we can expect that each zero of $d_{2}(B)$ in equation (3.8) is a limit of a zero of $c_{m+1}(B)$ as $m\to\infty$. We summarize the approximations of the zeros with respect to the expansions on the parameter $s(=2)$ and the numerical calculations of the zeros as the following table.
0th approx. 11st approx. 22nd approx. zero of $c_{30}(B)$
$k=0$ 0 11 $1.500000000$ $1.378489221$
$k=1$ 1-1 0 .4166666667-.4166666667 .2931662833-.2931662833
$k=2$ 4-4 3-3 3.033333333-3.033333333 3.035300946-3.035300946
$k=3$ 9-9 8-8 8.014285714-8.014285714 8.014303906-8.014303906
$k=4$ 16-16 15-15 15.00793651-15.00793651 15.00793924-15.00793924
$k=5$ 25-25 24-24 24.00505051-24.00505051 24.00505119-24.00505119
The 22nd order approximation is not good for small $k$, but it seems that the approximation would be better if $k$ is getting large.
@@
We can also expect that each zero of $c_{m+1}(B)$ converges as $m\to\infty$ for the case $s=2i$. We summarize the approximations of the zeros with respect to the expansions on the parameter $s(=2i)$ and the numerical calculations of the zeros as the following table.
0th approx. 11st approx. 22nd approx. zero of $c_{30}(B)$
$k=0$ 0 $i$ .5000000000+i-.5000000000+i .5406395812+.5331266960i-.5406395812+.5331266960i
$k=1$ 1-1 1+i-1+i .5833333333+i-.5833333333+i .5406395812+1.466873304i-.5406395812+1.466873304i
$k=2$ 4-4 4+i-4+i 3.966666667+i-3.966666667+i 3.968701175+i-3.968701175+i
$k=3$ 9-9 9+i-9+i 8.985714286+i-8.985714286+i 8.985730155+i-8.985730155+i
$k=4$ 16-16 16+i-16+i 15.99206349+i-15.99206349+i 15.99206621+i-15.99206621+i
$k=5$ 25-25 25+i-25+i 24.99494949+i-24.99494949+i 24.99495018+i-24.99495018+i
@@
The real parts of the first two numerical zeros for $c_{30}(B)$ coincide up to 10 digits, although the imaginary parts are $i$ for $k=0$ and $k=1$ on the 22nd order approximation.
@@
More than 20 zeros of $c_{40}(B)$ coincide with those of the zeros of $c_{30}(B)$ numerically up to 10 digits. We can also expect that each zero of $c_{m+1}(B)$ converges as $m\to\infty$ for the case $s=i$. We summarize the approximations of the zeros with respect to the expansions on the parameter $s(=i)$ and the numerical calculations of the zeros as the following table.
0th approx. 11st approx. 22nd approx. zero of $c_{30}(B)$
$k=0$ 0 .5i.5i .1250000000+.5i-.1250000000+.5i .1431861712+.5i-.1431861712+.5i
$k=1$ 1-1 1+.5i-1+.5i .8958333333+.5i-.8958333333+.5i .8775200792+.5i-.8775200792+.5i
$k=2$ 4-4 4+.5i-4+.5i 3.991666667+.5i-3.991666667+.5i 3.991792466+.5i-3.991792466+.5i
$k=3$ 9-9 9+.5i-9+.5i 8.996428571+.5i-8.996428571+.5i 8.996429618+.5i-8.996429618+.5i
$k=4$ 16-16 16+.5i-16+.5i 15.99801587+.5i-15.99801587+.5i 15.99801604+.5i-15.99801604+.5i
$k=5$ 25-25 25+.5i-25+.5i 24.99873737+.5i-24.99873737+.5i 24.99873742+.5i-24.99873742+.5i
By observing the zeros for two cases $s=2i$ and $s=i$, bifurcation structure is expected around $s=ai$ where $1<a<2$. In fact, if $s=i1.468768613785\cdots$, then two eigenvalues of the Mathieu equation coalesce and the value $s=i1.468768613785\cdots$ is called Mulholland-Goldstein double point. See [4] and references therein.
@@ 6.2. Whittaker-Hill equation
We describe a result in Section 4 on the singly confluent Heun equation. The recursive relation for the polynomial $c_{m+1}(B)$ was obtained in equations (3.13) and (3.14). It follows from Theorem 4.9 that the expansion of the zero of $c_{m+1}(B)$ is written as
@@
$\quad+\Bigl\{-\frac{1}{32}+\frac{P_{6}(k)}{(2k-3+\gamma+\delta)(2k-2+\gamma+\delta)^{3}(2k+\gamma+\delta)^{3}(2k+1+\gamma+\delta)}\Bigr\}s^{2}$
@@
We try to investigate the zeros of $c_{m+1}(B)$ numerically by setting the parameters to the particular values, and compare with the results by the approximation from the expansion in $s$. We restrict to the case
@@
Then, the expansion of the zero of $c_{m+1}(B)$ is written as
@@
We calculate the zeros of $c_{m+1}(B)$ numerically by fixing the values $s$ and $\alpha$.
We discuss the case $\alpha=5$ and s=1/100s=-1/100. The zeros of $c_{8}(B)$ is calculated numerically as
$-.02475005544,\;-1.025212444,\;-4.025020000,\;-9.025010357,$
@@
More than 20 zeros of $c_{40}(B)$ coincides with those of the zeros of $c_{30}(B)$ numerically up to 10 digits. Therefore, we can expect that each zero of $d_{2}(B)$ in equation (3.8) is a limit of a zero of $c_{m+1}(B)$ as $m\to\infty$.
We summarize the approximations of the zeros with respect to the expansions on the parameter s(=1/100)s(=-1/100) and the numerical calculations of the zeros as the following table.
0th approx. 11st approx. 22nd approx. zero of $c_{30}(B)$
$k=0$ 0 .0250000000-.0250000000 .0247500000-.0247500000 .02475005544-.02475005544
$k=1$ 1-1 1.025000000-1.025000000 1.025212500-1.025212500 1.025212444-1.025212444
$k=2$ 4-4 4.025000000-4.025000000 4.025020000-4.025020000 4.025020000-4.025020000
$k=3$ 9-9 9.025000000-9.025000000 9.025010357-9.025010357 9.025010357-9.025010357
$k=4$ 16-16 16.02500000-16.02500000 16.02500714-16.02500714 16.02500714-16.02500714
$k=5$ 25-25 25.02500000-25.02500000 25.02500568-25.02500568 25.02500568-25.02500568
@@
We discuss the case $\alpha=5$ and s=20s=-20. A part of the zeros of $c_{16}(B)$ is calculated numerically as
$2.051282315,\;31.28075111,\;54.63259197-14.18872668i,\;\dots,$
@@
$2.051136527,\;31.66623015,\;50.75984059,\;59.89047866-28.89267645i,\;\dots,$
and other zeros are not real numbers. We might expect that $c_{m+1}(B)$ has a zero around $2.051$ as $m\to\infty$, but it would not be true. It is shown by the numerical calculation that all zeros of $c_{50}(B)$ are not real numbers. A part of the zeros of $c_{89}(B)$ is calculated numerically as
@@
26 zeros of $c_{100}(B)$ are real numbers. We expect that the polynomial $c_{m+1}(B)$ has zeros around $11.72870190,\;-37.26280325,\;-59.78916204,\dots$, if $m$ is sufficiently large.